OpenStudy (anonymous):

Solve 3x-6y+5z=2 when x=2 and z = -2

OpenStudy (pfenn1):

Plug 2 in where x is and -2 for z and solve for y

OpenStudy (anonymous):

So substitute the values of x and z and solve for the remaining y..

OpenStudy (anonymous):

3(2)-6y+5(-2)=2

OpenStudy (anonymous):

The reason I ask this is because in the system 3x-6y+5z=2 3x+3y-z=5 5x+6y+5z=-6 When I try to plug in the values and solve for y, I get a strange answer... but when I plug it into the other equations, I get the correct answer. Why is this?

OpenStudy (anonymous):

I don't understand -6y = 8??

OpenStudy (anonymous):

WAIT

OpenStudy (anonymous):

Is it possible for it to be a fraction?

OpenStudy (anonymous):

NO FREAKING DUH... I SWEAR I'M HALF ASLEEP RIGHT NOW

OpenStudy (anonymous):

Oh ill wait:)

OpenStudy (anonymous):

4+2 = 6. lol

OpenStudy (anonymous):

lol, There you go!

OpenStudy (anonymous):

Im half asleep too!!

OpenStudy (anonymous):

lol :)

OpenStudy (anonymous):

Glad I could help...sorta! lol:)